# Uncertainty relations for overcomplete measurements from generalized equiangular tight frames

### v1

@@ III Main results
This section presents uncertainty relations for informationally overcomplete measurements generalized equiangular tight frames including conical $2$-designs. One shall address the entropic formulation as well as the case of probabilities for separate outcomes. All the derived relations will be exemplified.
@@ III.1 Entropic uncertainty relations
Informational entropies [33, 34] provide a flexible way to pose the uncertainty principle for finite-dimensional systems. In some respects, the entropic formulation have advantages in comparison with the traditional one [31]. Entropic uncertainty relations are important not only from the conceptual viewpoint, but also for a lot of applications [45, 46]. In particular, such relations are used in derivation of criteria to detect entanglement [36, 37] and steerability [38, 39, 40]. Statistics of the measurements of interest can be interpreted in two ways. First, it relates to a single generalized equiangular measurement. Second, one links to a set of generalized symmetric measurements. For these measurements [28], the involved sums of squared probabilities satisfy a certain relation. This fact leads to a lot of consequences. To avoid bulky expressions, we will use positive numbers
$\omega_{\mu}$ $=\frac{1}{\mathcal{C}}\>\frac{\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ ,$ (37)
$\mathcal{C}$ $=\sum_{\mu=1}^{M}\frac{\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ .$ (38)
@@
$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\,.$ (40)
@@
Proof. It was shown in appendix C of [28] that
$\mathrm{tr}(\boldsymbol{\rho}^{2})-\frac{1}{d}=\sum_{\mu=1}^{M}\frac{1}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\biggl{(}\sum_{j=1}^{N_{\mu}}\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})^{2}-\frac{a_{\mu}\gamma_{\mu}}{d}\biggr{)}.$ (42)
Combining this with (15) and (37) immediately gives
$\mathcal{C}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}=\mathrm{tr}(\boldsymbol{\rho}^{2})-\frac{1}{d}+\frac{1}{d}\,\sum_{\mu=1}^{M}\frac{\gamma_{\mu}}{a_{\mu}(b_{\mu}-c_{\mu})}\ ,$ (43)
where $I_{\mu}\equiv{I}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$. Using (39), we further write
@@
$\geq{L}_{\alpha}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}$ (45)
Here, the step (44) follows from (33) and the step (44) follows from convexity of the function $X\mapsto{L}_{\alpha}(X)$. Combining (43) with (45) completes the proof of (40).
For a conical $2$-design from collection of $M$ generalized equiangular lines, the paper [28] proved
@@
The statement of Proposition 1 poses Tsallis-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement. In particular, the inequality (40) is an entropic uncertainty relation with an average Tsallis $\alpha$-entropy. A distinction from the previous considerations is that averaging in the left-hand side of (40) is taken with generally unequal weights. This is a flip side of possibility to manipulate the coefficients of a generalized equiangular measurement. Another formulation of uncertainty relations uses the Rényi entropies.
@@
$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{\mathrm{tr}(\boldsymbol{\rho}^{2})\!\;d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,.$ (47)
@@
for $\alpha\in[1,2]$. In this interval, the inequality (50) leads to (47).
@@
The inequalities (47) and (48) are Rényi-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement. The first formula evaluates from below the Rényi entropy averaged over POVMs with the actual weights (37). In contrast to (48), it is proved only for $\alpha\in[1,2]$. The latter follows from convexity properties of the functions involved into consideration. The inequality (48) deals with a single generalized equiangular measurement. It holds for $\alpha\in(0,2]$ similarly to (41).
@@ III.2 Fine-grained uncertainty relations
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,.$ (52)
@@
$\varLambda_{p}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\leq\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N_{M}}\,\biggl{\{}1+\sqrt{N_{M}-1}\,\Bigl{(}N_{M}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+N_{M}(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,,$ (53)
@@
The statement of Proposition 3 imposes restrictions on maximal probabilities of the considered measurement. The result (52) estimates from above the maximal probability averaged over POVMs with the weights (37). In general, these weight are not equal. The second result (53) provides estimations of the maximal probability from below and above simultaneously. These results could be converted into inequalities with the corresponding min-entropies. We refrain from presenting the details here.
@@
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}\gamma_{\mu}^{-1}a_{\mu}\omega_{\mu}\biggr{)}\,.$ (57)
@@
$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N_{M}}\!\Bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\Bigr{)}\,,$ (58)
where the total number of outcomes $N_{M}$ is defined by (54).
Proof. By construction, the function $X\mapsto{F}_{N}(X)$ is concave. Due to (36), we therefore have
@@
The statement of Proposition 4 is an example of fine-grained uncertainty relations with the use of probabilities of two separate outcomes. It is not a surprise that the inequality (57) deals with averaging with the weights (37) generally unequal. In the paper [55], relations of such a kind were formulated for measurements assigned to a projective $2$-design. In particular, the case of $d+1$ MUBs was addressed assuming that $d$ is a prime power. Thus, Proposition 4 provides an extension of recent results to generalized equiangular measurements. It would be interesting to examine applications of these inequalities in quantum information, including schemes to detect entanglement and steerability.
@@ III.3 Examples
Let us illustrate the derived relations with some examples. It is instructive to begin with a qubit example. Following [28], we consider the operators
@@
To exemplify the case of conical $2$-design, we take the operators [28]
@@
$\gamma_{1}$ $=\frac{3\sqrt{5}-5}{4}\ ,\qquad\gamma_{1}=1-\gamma_{1}\,,$ (67)
@@
Dealing with a single measurement, one can rescale each of the used entropies by its maximal possible value. Then the curves all reduce to the same interval. This example includes five outcomes, so that we divide $H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ by $\ln_{\alpha}(5)$ and $R_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ by $\ln 5$. The rescaled values of the $\alpha$-entropies versus square of the Bloch vector are shown in Fig. 2 for $\alpha=0.8$ and the three directions. The lower bounds are calculated due to (41) and (48), respectively Similarly to the picture of Fig. 1, the deviation from the lower bounds is maximal for pure states. For the Bloch vector along the $z$-axis, it is around 9.1 % in the Tsallis case and 7.9 % in the Rényi case. In a relative scale, deviations are less than for the curves in Fig. 1. Indeed, this example deals with the number of outcomes that is larger.
@@
The authors of [26] showed how to build a complete set of $d+1$ mutually unbiased measurements in ${\mathcal{H}}_{d}$. In general, the parameter $\varkappa$ obeys $d^{-1}\leq\varkappa\leq 1$. The maximal efficiency $\varkappa=1$ corresponds to mutually unbiased bases (MUBs) with rank-one elements. In the considered example, the parameters (10)–(13) read as
@@
For the maximal efficiency $\varkappa=1$ corresponding to MUBs the inequality (77) reduces to one of the results proved in [55]. Then its right-hand side is strictly less than 1 for $d\geq 3$ and an impure state. For $\varkappa<1$ and $d\geq 3$, the inequality (77) is nontrivial for all states. Using MUMs, a generalized equiangular measurement can be built due to (15). We refrain from presenting the details here.

### v3

@@ III Main results
This section presents uncertainty relations for quantum measurements from generalized equiangular tight frames including conical $2$-designs. One shall address the entropic formulation as well as the case of probabilities of separate outcomes. All the derived relations will be exemplified.
@@ III.1 Entropic uncertainty relations
Informational entropies [34, 35] provide a flexible way to pose the uncertainty principle for finite-dimensional systems. In some respects, the entropic formulation have advantages in comparison with the traditional one [32]. Entropic uncertainty relations are important not only from the conceptual viewpoint, but also for a lot of applications [46, 47]. In particular, such relations are used in derivation of criteria to detect entanglement [37, 38] and steerability [39, 40, 41]. Statistics of the measurements of interest can be interpreted in two ways. First, it deals with a single generalized equiangular measurement. Second, one links to a set of generalized symmetric measurements. For these measurements [29], the involved sums of squared probabilities satisfy a certain relation. This fact leads to a lot of consequences. To avoid bulky expressions, we will use the positive numbers
$\omega_{\mu}$ $=\frac{\mathcal{C}^{-1}}{x_{\mu}-y_{\mu}}=\frac{\mathcal{C}^{-1}\gamma_{\mu}^{2}}{a_{\mu}^{2}(b_{\mu}-c_{\mu})}\ ,$ (34)
$\mathcal{C}$ $=\sum_{\mu=1}^{M}\frac{1}{x_{\mu}-y_{\mu}}\ .$ (35)
@@
$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,.$ (37)
@@
Proof. To prove (37), we denote (84) by $I_{\mu}$ and further write
@@
$\geq{L}_{\alpha}\biggl{(}\,\sum_{\mu=1}^{M}\omega_{\mu}\,I_{\mu}\biggr{)}\,.$ (40)
Here, the step (39) follows from (30) and the step (40) follows from convexity of the function $X\mapsto{L}_{\alpha}(X)$. Combining (85) with (40) completes the proof of (37), since the function $X\mapsto{L}_{\alpha}(X)$ decreases.
For a conical $2$-design from collection of $M$ generalized equiangular lines, the author of [29] proved that
@@
The statement of Proposition 1 gives Tsallis-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement. In particular, the inequality (37) is an entropic uncertainty relation with average Tsallis $\alpha$-entropy. A distinction from the previous considerations is that averaging in the left-hand side of (37) is taken with generally unequal weights. This is a flip side of possibility to manipulate the coefficients of a generalized equiangular measurement. In contrast to (37), the inequality (38) deals with a single measurement. Instead of (85), the corresponding index of coincidence is calculated exactly due to (41). The weights $\omega_{\mu}$ in the left-hand side of (37) are determined by the coefficients of a generalized equiangular measurement. Since $\mathrm{tr}(\boldsymbol{\rho}^{2})=1$ for a pure state, the inequalities (37) and (38) then give
$\sum_{\mu=1}^{M}\omega_{\mu}\,H_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $\geq{L}_{\alpha}\biggl{(}\frac{d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,,$ (42)
$H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ $\geq{L}_{\alpha}\bigl{(}S+(\sigma-S)d^{-1}\bigr{)}\,.$ (43)
These relations actually hold for all states, because the function $X\mapsto{L}_{\alpha}(X)$ decreases. Such inequalities are typically used in criteria to detect nonclassical correlations. Together with the weights $\omega_{\mu}$, the right-hand side of (42) can be treated as an estimation of Tsallis entropies related to $M$ generalized symmetric POVMs. In a similar vein, the right-hand side of (43) is an entropic characteristic of the given conical $2$-design. The next formulation of uncertainty relations uses the Rényi entropies.
@@
$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{\mathrm{tr}(\boldsymbol{\rho}^{2})\!\;d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,.$ (44)
@@
for $\alpha\in[1,2]$. In this range, the inequalities (47) and (85) lead to (44), since the function (48) decreases.
@@
The inequalities (44) and (45) are Rényi-entropy uncertainty relations for measurements assigned to a generalized equiangular measurement. The first formula estimates from below the Rényi entropy averaged over POVMs with the actual weights (34). In contrast to (45), it is proved only for $\alpha\in[1,2]$. The latter follows from convexity properties of the functions involved into consideration. The inequality (45) deals with a single generalized equiangular measurement. It holds for $\alpha\in(0,2]$ similarly to (38). For a pure state, the inequalities (44) and (45) give
$\sum_{\mu=1}^{M}\omega_{\mu}\,R_{\alpha}({\mathcal{E}}_{\mu};\boldsymbol{\rho})$ $\geq\frac{1}{1-\alpha}\,\ln\biggl{\{}1+(1-\alpha)L_{\alpha}\biggl{(}\frac{d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\biggr{\}}\,,$ (49)
$R_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ $\geq\frac{1}{1-\alpha}\,\ln\Bigl{\{}1+(1-\alpha)L_{\alpha}\bigl{(}S+(\sigma-S)d^{-1}\bigr{)}\Bigr{\}}\,.$ (50)
Similarly to (42) and (43), these relations hold for all states and have potential applications. The right-hand sides of (49) and (50) can be interpreted as an estimate of Rényi entropies assigned to the corresponding measurements.
@@ III.2 Fine-grained uncertainty relations
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,.$ (51)
@@
$\varLambda_{p}\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\leq\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})\leq\frac{1}{K}\,\biggl{\{}1+\sqrt{K-1}\,\Bigl{(}KS\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+K(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,,$ (52)
@@
The statement of Proposition 3 imposes restrictions on maximal probabilities of the considered measurement. The result (51) estimates from above the maximal probability averaged over POVMs with the weights (34). In general, these weight are not equal. The second result (52) provides estimations of the maximal probability from below and above simultaneously. These results could be converted into inequalities with the corresponding min-entropies. We refrain from presenting the details here. The state-independent counterparts of (51) and (52) respectively read as
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{1\leq{j}\leq{N}}{\max}\,\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})$ $\leq\frac{1}{N}\,\biggl{\{}1+\sqrt{N-1}\,\biggl{(}\frac{Nd-N}{\mathcal{C}{d}}+\frac{N}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}-1\biggr{)}^{\!1/2}\,\biggr{\}}\,,$ (55)
$\underset{\mu,j}{\max}\,\mathrm{tr}({\mathsf{P}}_{\mu,j}\boldsymbol{\rho})$ $\leq\frac{1}{K}\,\biggl{\{}1+\sqrt{K-1}\,\Bigl{(}KS+K(\sigma-S)d^{-1}-1\Bigr{)}^{\!1/2}\,\biggr{\}}\,.$ (56)
These inequalities are obtained by replacing $\mathrm{tr}(\boldsymbol{\rho}^{2})$ with its maximal value one. Note that only an estimate from above appears in (56). The formulas (55) and (56) allow us to estimate the maximal probabilities from above in terms of the characteristics of POVMs solely.
@@
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,.$ (58)
@@
$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{K}\!\bigl{(}S\!\;\mathrm{tr}(\boldsymbol{\rho}^{2})+(\sigma-S)d^{-1}\bigr{)}\,,$ (59)
where the total number of outcomes $K$ is defined by (53).
Proof. By construction, the function $X\mapsto{F}_{N}(X)$ is increasing and concave. Due to (33), we therefore have
@@
The statement of Proposition 4 is an example of fine-grained uncertainty relations with the use of probabilities of two separate outcomes. It is not a surprise that the inequality (58) realizes averaging with the weights (34) generally unequal. In the paper [56], similar relations with equal weights were formulated for measurements assigned to a projective $2$-design. In particular, the case of $d+1$ MUBs was addressed assuming that $d$ is a prime power. Thus, Proposition 4 provides an extension of recent results to generalized equiangular measurements. To convert (58) and (59) into the state-independent form, we replace $\mathrm{tr}(\boldsymbol{\rho}^{2})$ with its maximal value one, whence
$\sum_{\mu=1}^{M}\omega_{\mu}\,\underset{i\neq{j}}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{E}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{E}}_{\mu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{N}\biggl{(}\frac{d-1}{\mathcal{C}{d}}+\frac{1}{d}\,\sum_{\mu=1}^{M}w_{\mu}\omega_{\mu}\biggr{)}\,,$ (61)
$\underset{(\mu,i)\neq(\nu,j)}{\max}\,\bigl{\{}\mathrm{tr}({\mathsf{P}}_{\mu,i}\boldsymbol{\rho})+\mathrm{tr}({\mathsf{P}}_{\nu,j}\boldsymbol{\rho})\bigr{\}}\leq{F}_{K}\!\bigl{(}S+(\sigma-S)d^{-1}\bigr{)}\,.$ (62)
It would be interesting to examine applications of these inequalities in quantum information, including schemes to detect entanglement and steerability.
@@ III.3 Examples
Let us illustrate the derived relations with some examples. It is instructive to begin with a qubit. Its density matrix reads as
[…]
where $r_{x}=\sin\theta\cos\varphi$, $r_{y}=\sin\theta\sin\varphi$ and $r_{z}=\cos\theta$ are the components of the Bloch vector. By $\theta$ and $\varphi$, we mean usual angles on the Bloch sphere. Three eigenbases of the Pauli matrices $\boldsymbol{\tau}_{x}$, $\boldsymbol{\tau}_{y}$ and $\boldsymbol{\tau}_{z}$ are mutually unbiased. This case deals with three von Neumann measurements. Here, the weights (34) are equal to one third. The six states of the three eigenbases also form a $3$-design [64] and, herewith, a $2$-design. The resulting POVM consists of elements of the form $3^{-1}|b_{\mu,j}\rangle\langle{b}_{\mu,j}|$ with $\mu=x,y,z$ and $j=\pm$. It follows from (37) and (38) that, for $\alpha\in[0,2]$,
$\frac{1}{3}\sum_{\mu=x,y,z}H_{\alpha}({\mathcal{B}}_{\mu};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{3+r^{2}}{6}\biggr{)}\,,$ (64)
$H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})\geq{L}_{\alpha}\biggl{(}\frac{3+r^{2}}{18}\biggr{)}\,,$ (65)
where $r^{2}=r_{x}^{2}+r_{y}^{2}+r_{z}^{2}$ is square of the Bloch vector. The uncertainty relations (64) and (65) were actually reported in [56]. A utility of these inequalities was also discussed therein.
The next example deals with unequal weights (34). Following [29], we consider the operators
@@
To exemplify the case of conical $2$-design, we use the operators [29]
@@
$\gamma_{1}$ $=\frac{3\sqrt{5}-5}{4}\ ,\qquad\gamma_{2}=1-\gamma_{1}\,,$ (73)
@@
Dealing with a single measurement, one can rescale each of the used entropies by its maximal possible value. Then the curves all reduce to the same interval. This example includes five outcomes, so that we divide $H_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ by $\ln_{\alpha}(5)$ and $R_{\alpha}({\mathcal{P}};\boldsymbol{\rho})$ by $\ln 5$. The rescaled values of the $\alpha$-entropies versus square of the Bloch vector are shown in Fig. 2 for $\alpha=0.8$ and the three directions. The lower bounds are calculated due to (38) and (45), respectively. Similarly to the picture of Fig. 1, the deviation from the lower bounds is maximal for pure states. For the Bloch vector along the $z$-axis, it is around 9.1 % in the Tsallis case and 7.9 % in the Rényi case. In a relative scale, deviations are less than for the curves in Fig. 1. Indeed, this example deals with the number of outcomes that is larger.
@@
The authors of [27] showed how to build a complete set of $d+1$ mutually unbiased measurements in ${\mathcal{H}}_{d}$. In general, the parameter $\varkappa$ obeys $d^{-1}\leq\varkappa\leq 1$. The maximal efficiency $\varkappa=1$ corresponds to mutually unbiased bases (MUBs). In the considered example, the parameters (9)–(12) read as
@@
For the maximal efficiency $\varkappa=1$ corresponding to MUBs the inequality (83) reduces to one of the results proved in [56]. Then its right-hand side is strictly less than 1 for $d\geq 3$ and a state that is not pure. For $\varkappa<1$ and $d\geq 3$, the inequality (83) is nontrivial for all states. Using MUMs, a generalized equiangular measurement can be built due to (14). We refrain from presenting the details here.